

********************************************************************
UARS WEEKLY STATUS REPORT - 12 JAN 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible. We
continue to gather enough data to meet all the SOLSTICE science objectives
during this period of reduced tape recorder operations.
This week we successfully executed a number of observations of Mercury. A
special thank you to Rich Beck and the folks at the MPG for their help in
scheduling this activity.
***********
SUSIM
***********
SUSIM performed its daily solar scans of the sun during periods of
predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime
winds when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG assisted with the scheduling of observations of
Mercury by the SOLSTICE instrument.
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 9 February 2001.
***********
CDHF
***********
UARS CDHF REHOST:
- We shall be sending out to all PI/SW groups, responses to all
concerns (SUSIM,GPI), we have had regarding the results of comparing
rehost and current system results. We have found reasonable reason for
differences and we will send this information out in the next few days.
- Our plans are to go operational on the rehost system on January 31,2001.
Operational status means that L0,NRT,Quicklook and correlative ingests
and current day processing will be done on the REHOST system. The
old system will still be available until we migrate all old UARS data
to the REHOST system.
UARS CDHF
- The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
- ***LEVEL0 for UARS DAYS 3399 and 3403 RE-INGESTED for "Better Quality"***
1 day ATT (7.3) PROD (3403) (1R-1C=0L)
1 day OBC (3.4) PROD (3403) (1R-1C=0L)
1 day SEN (5.0) PROD (3399) (1R-1C=0L)
- The CDHF is "catching up" with the HALOE production processing. LEVELS 2-3
has been on hold since 21-December-2000. The Auxiliary (Sol) data was
delayed due to extensive preventive maintenance at the Solar Terrestrial
Dispatch System. The backlog of data started becoming available 8-Jan-2001.
HALOE software representative has been notified.
DAAC DATA:
- A priority is being given to the preview and loading of MLS V5 into the
DAAC Transfer Tables. The data created in the month of October and
November, 2000 has been loaded into the DAAC Transfer Tables. Plans are
to start previewing the data created in the month of December, 2000.
- Data created in the month of October, 2000 has been loaded into the
DAAC Transfer Tables for HALOE, NMC, and UKMO. DAAC has been notified.
***********
FOT
***********
REPORTING PERIOD
This report covers 30 December 2000 (Orbit 50851, GMT Day 365) through 05
January 2001 (Orbit 50955, GMT Day 005).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 365/2245 50865 Power OFF
PEM 365/2245 50865 Power OFF
SUSIM 365/2245 50865 Power OFF
HRDI 366/0029 50866 Low Power mode
ACRIM 002/0004 50895 Power ON
SOLSTICE 002/0143 50896 Power ON
SUSIM 002/0324 50898 Power ON
PEM 002/1244 50903 ZEPS/HEPS Day only
HALOE 004/0006 50925 Power ON
A Forward-to-Reverse Yaw Around maneuver (number 95) was successfully executed
in Orbit 50878 beginning 366/14:57:50 and completing at 366/15:34:50 on 31
December 2000. The SOLSTICE instrument was in day only and all other
instruments were in their off or standby mode (including HRDI and PEM). Due to
the SA position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 50.57 minutes. Post maneuver analysis of
data indicate that during the Yaw Around, the end-of-night load bus voltage was
25.14 volts and the average depths of discharge (DODs) on Batteries 2 and 3
were approximately 22.51and 21.90 percent respectively. The next (reverse-to-
forward) YAW is scheduled for 09 February 2001 (DOY 040).
The beta angle ranged between the angles of 05.3 degrees to 30.4 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 77 degrees on 21 January 2001 (DOY 021).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge realy. Battery 2 and Baterry 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 and 1.0 degrees C respectively. Additional battery performance data
for each battery on 28 December 2000 and 05 January 2001 are provided in
Appendix D.
The clock error ranged from ~10.3 to 12.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.
A summary is listed below and the details of ALERT and Out-Of-Limits
occurrences are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
*Data not available this week.
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/315 - 2287.496493
· Transponder B: 00/314 - 2287.498584
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/342 show the observatory data loss to be 300 hours, 08
minutes, 41.256 seconds. This shows an overall increase of 6 minutes and 52.288
seconds since the last report period. This is a 0.39 percent data loss, which
equals a 99.61 percent data capture for the mission.
The increase of 6 minutes and 52.288 seconds occurred over one (5) days:
· 01 minutes and 16.752 seconds on Day 364 due to a data hit.
· 01 minutes and 18.656 seconds on Day 365 due to a data hit.
· 13.552 seconds on Day 366 due to a data hit.
· 03 minutes and 36.032 seconds on Day 1 due to a command uplink failure
TTR#23380.
· 27.296 seconds on Day 2 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Three Axis Magnetometer (TAM) was put back into update mode on 03 January
2001. After tuning some of the noise parameters and sensor alignments, the TAM
is performing well enough to keep it in update mode continuously. The FOT
plans to fly on TAM only through the weekend and into next week, contingent on
the FSS residuals not exceeding 1500 arc seconds.
The Flight Software patch to account for the nonorthoganality of the TAM was
tested this week and is scheduled to be uplinked on 09 January 2001. This
patch should further improve the TAM's performance. If the TAM continues to
operate this well, UARS will be able to fly through the upcoming high beta
period on TAM alone without having to perform a yaw offset maneuver.
MEETINGS
The FOT attended a TAM Attitude Meeting on 02 January 2001.
The FOT attended a UARS Monthly Status Meeting on 03 January 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for December 31, 2000 (DOY 366). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 January 2001
*Data not available this week.
_____________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 January 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 January 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period
___________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 January 2001
28 Dec 2000 - GMT Day 363 Beta = 10.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------ ------ ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +11.2/-22.4 +4.51/+3.87 +29.6/-13.2 26.2 1.020 19.9
3 +22.4/-22.4 +3.23/+2.60 +28.8/-12.8 26.1 1.020 19.3
05 Jan 2001 - GMT Day 005 Beta = 30.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------ ------ ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +0.0/-44.8 +5.47/+4.51 +16.0/-15.2 25.9 1.030 21.0
3 +11.2/-22.4 +4.51/+3.55 +15.6/-15.2 25.8 1.030 20.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 January 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-000 FOT 50890 001/1605 23380 System failover at WSGT
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-047 FOT see list Nov 00 14170 Generic Late Acq.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
01-000 FOT 50890 001/1605 23380 System failover at WSGT
SEVNTFW01/01